OpenStudy (anonymous):

evaluate the following integral. int (2xe^{x}^{2})/(1+e^{x}^{2})

OpenStudy (anonymous):

\[\int\limits_{}^{} 2xe ^{x ^{2}}/1+e ^{x ^{2}}\]

OpenStudy (amistre64):

the top is the derivative of the bottom right?

OpenStudy (anonymous):

this is what I have

OpenStudy (amistre64):

\[\int\frac{dx}{x}=ln|x|+C\]

OpenStudy (anonymous):

\[\int\limits_{}^{} 2xe ^{x ^{2}}/1+e ^{x ^{2}}dx\] \[\int\limits_{}^{} 2e ^{x ^{2}}x/1+e ^{x ^{2}}\] \[2\int\limits_{}^{} e ^{x ^{2}}x/1+e ^{x ^{2}}\] \[u=e ^{x ^{2}}\] \[du=2e ^{x ^{2}}\] \[\int\limits_{}^{} 1/u\] \[\ln(u)+c\] \[\ln(e ^{x ^{2}}+1)+c\]

OpenStudy (anonymous):

the problem i am having trouble with is the step right before i integrate \[2\int\limits_{}^{}x^{x ^{2}}x/1+e ^{x ^{2}}\] does the top portion of the equation equal 1?????

OpenStudy (amistre64):

in the case above; u = e^(x^2); du = 2x e^(x^2)

OpenStudy (anonymous):

that first x is supposed to be an e

OpenStudy (amistre64):

why are you pulling out the 2?

OpenStudy (anonymous):

because the 2 is a constant

OpenStudy (amistre64):

just becasue it is a constant doent not mean you pull it out; it is very useful to us so leave it in

OpenStudy (anonymous):

so i was under the impression to always pull out the constant. how do i know when and when not to pull out the constant?

OpenStudy (amistre64):

du \(2x\ e^{x^2}\) --- = ------- u \(1+e^{x^2}\) you only pull out a constant if its useless to the over all integration

OpenStudy (amistre64):

the 2 helps to establish the fact that the top is the derivative of the bottom

OpenStudy (amistre64):

we know how to integrate du/u .... so ues it :)

OpenStudy (anonymous):

i see. that helps clear things up. thanks

OpenStudy (amistre64):

youre welcome